OpenStudy (anonymous):

What is the simplified form of the square root of 72n^13 ?

OpenStudy (anonymous):

\[\large \sqrt{72 \times n^{13}} \implies \sqrt{2 \times 2 \times 2 \times 3 \times 3 \times n^6 \times n^6 \times n}\] \[\large \sqrt{72n^{13}} = 2 \times 3 \times n^6 \sqrt{2n} = ??\]

OpenStudy (anonymous):